# Lightweight and Interpretable Transformer via Mixed Graph Algorithm Unrolling for Traffic Forecast

### v2

@@ F.1 ADMM Blocks Setup
We initialize $\mu_{u},\mu_{d,1},\mu_{d,2}$ with 3, and the $\rho,\rho_{u},\rho_{d}$ with $\sqrt{N/(T+S+1)}$, where $N$ is the number of sensors and $T+S+1$ is the full length of the recovered signal (both the observed part and the part to be predicted). Those parameters are trained every ADMM layer. We initialize the CGD parameters ($\bm{\alpha}$}, $\bm{\beta}$})s as 0.08, and tune them for every CGD iteration. We clamp each element of the CGD step size $\bm{\alpha}$} to [0, 0.8] for stability. Each element of the momentum term coefficient $\bm{\beta}$} is clamped to be non-negative.
@@ F.3 Feature Extractors Setup
Input Laplacian embeddings Position embeddings Concatenate embedded input
@@ F.3.1 Input Embeddings
We generate non-parametric spatial and temporal embeddings for each node in our mixed spatio-temporal graph, see Figure 4(a). For spatial embeddings, we first generate the adjacency matrix ${\mathbf{W}}^{s}$ for the real-world connection of monitor stations using the physical distances $\text{d}^{S}(i,j)$ with
$w_{i,j}^{S}=w_{j,i}^{S}=\exp\left[-\frac{(\text{d}^{S}(i,j))^{2}}{50\times\text{Var}(\text{d}^{S})}\right],$ (66)
where $\text{Var}(\text{d}^{S})$ denotes the variance of the physical distances. We use Laplacian embedding supposed by Dwivedi & Bresson (2020), which uses the eigenvectors of the $k^{S}$-smallest non-zero eigenvalues of the Laplacian matrix ${\mathbf{L}}^{s}={\mathbf{D}}^{s}-{\mathbf{W}}^{s}$ to generate $k$-dimensional embeddings. Our experiments set $k^{S}=5$.
For temporal embeddings, we use the sine-cosine position embeddings (Vaswani et al., 2018):
@@
$e[t,2i+1]=$ $\cos(t/10000^{i}),i=0,1,\dots,\left\lfloor\frac{k^{T}}{2}\right\rfloor-1.$ (67)
We follow Feng & Tassiulas (2022) to use real-time-stamp, time-in-day and day-in-week embeddings of dimensions 10, 6, 4.
@@ F.3.2 Feature Extraction from embeddings
We use a variant of GraphSAGE (Hamilton et al., 2017) to aggregate spatial features and introduce Temporal History Aggregation (THA) to aggregate spatial features. The GraphSAGE module aggregates the inputs of each node’s $k$-NNs with a simple and uniformed linear layer to generate $K$-dimension spatial features for each node of each time step. The THA module aggregates the embedded inputs of past $W$ time steps for each node with a uniformed linear layer, where $W$ is the size of the time window. For both GraphSAGE and THA modules, we pad zeros to inputs as placeholders, and use Swish (Ramachandran et al., 2017) as acitvation functions with $\beta=0.8$.
@@ F.5 Training Setup
Dataset PeMS03 PeMS04 METR-LA
@@
Batch size 12 12 16
Appendix G Training Performance w.r.t. Training Data Cost
@@
Appendix H Ablation Studies
H.1 Ablation Studies for $\ell_{2}$- and $\ell_{1}$- norms in Direct Graph Modeling
First, we discuss the effectiveness of the new directed graph priors, DGLR and DGTV, as well as the benefit of directed-graph modeling of temporal relations. We separately test the effects of DGLR and DGTV by removing the prior and unrolling the ADMM algorithm. We denote these models as “w/o DGTV” and “w/o DGLR”. To test the benefit of the directed-graph modeling, we change the temporal graph ${\mathcal{G}}^{d}$ into a fully undirected graph ${\mathcal{G}}^{n}$ while maintaining the connections, and replace the directed graph priors with a GLR prior ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ with the same coefficient $\mu_{n}=\mu_{d,2}$. We denote the undirected-time unrolling model as “UT”. Detailed derivations of the ADMM algorithm for all the above ablation studies are in Appendix I.
Table 5: Performance comparison in RMSE / MAE / MAPE(%) of the ablation studies of directed-graph modeling and directed graph priors in 60-minute forecasting with PEMS03 and METR-LA.
Dataset Mixed-Graph UT w/o DGTV w/o DGLR
PEMS03 27.67 / 17.46 / 17.72 27.52 / 17.87 / 18.82 27.78 / 17.85 / 17.90 30.89 / 20.02 / 21.10
METR-LA 12.34 / 5.18 / 11.80 12.51 / 5.42 / 12.11 12.36 / 5.40 / 12.31 12.41 / 5.35 / 12.20
H.2 Sensitivity of $k$, $W$ in Graph Construction
@@
 4*  6* 27.67 17.46 17.72  6*  6* 12.37 5.27 11.91
4 4 29.54 17.09 17.92 4 6 12.43 5.24 11.87
4 8 27.45 17.25 18.35 8 6 12.61 5.30 11.92
6 6 28.98 17.33 18.23 6 4 12.44 5.40 11.85
8 6 45.37 18.15 20.59 6 8 12.44 5.35 11.96
H.3 Sensitivity of Number of Blocks and Layers in the Unrolling Model
@@
 5*  25* 27.67 17.46 17.72 12.34 5.18 11.80
5 20 27.95 17.32 18.02 12.53 5.35 12.15
5 30 27.86 17.73 18.80 12.54 5.31 11.94
4 25 27.77 17.63 18.61 12.45 5.33 12.00
6 25 27.84 17.54 18.61 12.37 5.27 11.90
H.4 Ablation Study for the Splitting-$\ell_{2}$-term Strategy
We also ran ablations of the splitting-term strategy by optimizing ${\mathbf{x}}^{\tau+1}$ using only eq. 7. CG in solving eq. 7 converges much slower than in solving eq. 10-eq. 12, and the unrolled model can become numerically unstable in training, indicating that splitting the $\ell_{2}$ terms can stabilize the training process.
@@ I.1 Optimization & Algorithm without DGTV Term
By removing the DGTV term $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ in eq. 4, and introducing the auxiliary variables ${\mathbf{z}}_{u},{\mathbf{z}}_{d}$ as eq. 9, we rewrite the unconstrained version of the optimization by augmented Lagrangian method as
$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{z}}_{d}+({\bm{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$
$+({\bm{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|_{2}^{2}.$ (68)
@@
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\bm{\gamma}}_{u}^{\tau}+{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau},$ (69)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (70)
$\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{z}}_{d}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}.$ (71)
The updating of ${\bm{\gamma}}_{u}^{\tau},{\bm{\gamma}}_{d}^{\tau}$ follows eq. 18.
@@ I.2 Optimization & Algorithm without DGLR Term
By removing the DGLR term ${\mathbf{x}}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ in eq. 4 and introduce only the auxiliary variables $\boldsymbol{\phi}$} in eq. 5 and ${\mathbf{z}}_{u}$ in eq. 9, we rewrite the unconstrained version of the optimization of ${\mathbf{x}}^{\tau+1}$ by augmented Lagrangian method as
$\min_{x,{\mathbf{z}}_{u}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{2}^{2}+({\bm{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$ (72)
@@
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}$ $={\mathbf{H}}^{\top}{\mathbf{y}}+({\mathbf{L}}^{d}_{r})^{\top}(\frac{\gamma^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}$ (73)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}$ $=\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}$ (74)
The optimization of ${\boldsymbol{\phi}}^{\tau+1}$ is the same as eq. 15, and the solution is the same as eq. 16. The updating of ${\bm{\gamma}},{\bm{\gamma}}_{u}$ follows eq. 17 and eq. 18.

### v3

@@ F.1 ADMM Blocks Setup
We initialize $\mu_{u},\mu_{d,1},\mu_{d,2}$ with 3, and the $\rho,\rho_{u},\rho_{d}$ with $\sqrt{N/(T+S+1)}$, where $N$ is the number of sensors and $T+S+1$ is the full length of the recovered signal (both the observed part and the part to be predicted). Those parameters are trained every ADMM layer. We initialize the CGD parameters ($\boldsymbol{\alpha}$}, $\boldsymbol{\beta}$})s as 0.08, and tune them for every CGD iteration. We clamp each element of the CGD step size $\boldsymbol{\alpha}$} to [0, 0.8] for stability. Each element of the momentum term coefficient $\boldsymbol{\beta}$} is clamped to be non-negative.
@@ F.3 Feature Extractors Setup
Input Spatial embeddings Temporal embeddings Concatenate embedded input
@@ F.3.1 Input Embeddings
We generate low-parametric spatial and temporal embeddings for each node in our mixed spatio-temporal graph, see Figure 7(a). For spatial embeddings, we embed each monitoring station with a 5-dimensional learnable vector, and broadcast the embeddings to every time step in the product graph. For temporal embeddings, we follow feng2022 to use real-time-stamp, time-in-day and day-in-week embeddings. To achieve better fitting capacity while keeping the model size small, we set only the time-in-day and day-in-week embeddings learnable, and use fixed embeddings for real time stamps. We use a fixed 10-dimensional sine-cosine position embeddings (tensor2tensor) for each time stamp:
@@
$e[t,2i+1]=$ $\cos(t/10000^{i}),i=0,1,2,3,4,$ (69)
and set dimensions of learnable time-in-day and day-in-week embeddings as 6 and 4. All embeddings are shared across all graph learning modules.
@@ F.3.2 Feature Extraction from embeddings
We use a variant of GraphSAGE (hamilton2017inductive) to aggregate spatial features and introduce Temporal History Aggregation (THA) to aggregate spatial features. The GraphSAGE module aggregates the inputs of each node’s $k$-NNs with a simple and uniformed linear layer to generate $K$-dimension spatial features for each node of each time step. The THA module aggregates the embedded inputs of past $W$ time steps for each node with a uniformed linear layer, where $W$ is the size of the time window. For both GraphSAGE and THA modules, we pad zeros to inputs as placeholders, and use Swish (ramachandran2017swish) as acitvation functions with $\beta=0.8$.
@@ F.5 Training Setup
Dataset PeMS03 PeMS08 METR-LA
@@
Batch size 12 16 16
Appendix G Further Experiments
G.1 Additional Experimental Results
Table 5 shows the performance comparison between our model and baselines in 30-/60-minute forecast on the PEMS08 (9346058) dataset which contains 170 nodes and 295 edges. Results show that our model also ranks top-3 in at least one metric of all experiments, and even gives the smallest MAPE among all models.
Table 5: Experimental results on additional dataset PEMS08 for 30- and 60-minute forecast.The indication of bolded, highlighted and underlined entries is the same as Table 1.
Horizons 30 minutes (6 steps) 60 minutes (12 steps)
Model RMSE MAE MAPE (%) RMSE MAE MAPE (%)
VAR 26.81 17.74 12.69 28.59 18.99 13.52
STGCN 36.58 24.39 16.83 50.61 33.01 20.57
STSGCN 26.73 17.43 11.48 30.22 19.74 12.90
GMAN 24.84 15.73 10.87 26.22 16.83 13.64
ST-Wave 26.85 17.19 11.08 27.28 16.85 11.77
PDFormer 22.92 13.91 9.17 27.36 17.61 12.90
STAEFormer 26.54 16.62 10.88 32.68 20.78 14.86
Graph WaveNet 30.95 21.65 13.56 40.24 29.12 17.79
AGCRN 34.31 19.91 10.88 38.09 22.78 12.75
SITD 27.40 18.11 12.10 33.12 22.66 20.08
SimpleTM 24.44 15.07 9.09 32.48 20.26 12.34
Ours 25.92 16.22 10.17 27.78 17.22 11.11
G.2 Experiments on the Entire Datasets
To further strengthen our proposed method, we ran the 60-minute forecast experiment on the PEMS03 dataset without sampling with 70 epochs of training, and the results are shown in Table 6. Our unrolled network still ranks top-3 in at least one of the metrics for each task, indicating the consistency of our performance with subsampling.
Table 6: Performance comparison of our model to the baselines on a 60-minute forecast for the entire PEMS03 dataset. The indication of bolded, highlighted and underlined entries is the same as Table 1.
Dataset PEMS03 METR-LA
Model RMSE MAE MAPE (%) RMSE MAE MAPE (%)
VAR 30.53 18.30 19.70 12.57 7.05 13.48
STGCN 36.77 21.81 19.37 12.27 5.88 11.79
STSGCN 28.17 17.46 16.86 12.73 5.11 10.66
PDFormer 25.05 14.78 15.53 12.32 4.77 10.20
GMAN 28.44 17.04 22.82 13.75 6.44 12.72
GWN 28.68 17.08 16.49 13.82 6.41 11.92
AGCRN 27.65 15.77 15.02 12.18 4.74 10.50
STID 26.80 15.14 16.45 12.00 4.63 10.14
SimpleTM 24.51 15.32 15.55 12.86 5.67 9.89
Ours 26.71 16.36 17.18 12.22 5.35 11.87
G.3 Inference-time Computational and Memory Cost
We compute the inference-time computational and memory cost on 60-minute forecasting for the PEMS03 dataset. We calculate the total floating-point operations (FLOPs) during inference as the computational cost with a batch size of 1, and evaluate the peak allocated memory by CUDA for each batch of size 32 as the memory cost. Results in Table 7 show that our unrolled network has the least computation cost among all the baselines except AGCRN and STID, which are generally consistently outperformed by our model. Our model’s inference-time memory cost is only larger than SimpleTM and the two baselines mentioned above. In conclusion, our lightweight transformer reduces the computational cost to only 4.9% of transformer-based PDFormer and 11.2% of the GAT-based GMAN with reduced or at least comparable memory cost.
Table 7: Comparison of computational cost (GFLOPs) for each data point and inference memory cost (GB) for a batch of size 32 in 60-minute forecasting for the PEMS03 dataset.
Model Ours STGCN GMAN ST-Wave PDFormer
Computation (GFLOPS) 0.087 0.013 $\times 12^{\dagger}$ 0.777 4.496 1.771
[…]
Model STAEFormer GWN AGCRN STID SimpleTM
[…]
    The computational and memory cost for STSGCN is not included due to the limitation of applying tools such as thop in the public code.
  • $\dagger$
    The $\times 12$ notation in the table means that the model predicts the outputs recursively.
G.4 Training Performance w.r.t. Training Data Cost
@@ G.3 Inference-time Computational and Memory Cost
Appendix H Further Ablation Studies for Hyperparameter Sensitivity
H.1 Sensitivity of $k$, $W$ in Graph Construction
@@
 4*  6* 26.96 16.58 18.07  6*  6* 12.17 5.27 11.78
4 4 27.23 16.80 17.73 4 6 12.28 5.24 11.86
4 8 27.30 17.34 19.34 8 6 12.45 5.26 11.91
6 6 29.57 18.00 19.43 6 4 12.38 5.23 11.84
8 6 28.40 17.99 18.48 6 8 12.52 5.42 12.29
H.2 Sensitivity of Number of Blocks and Layers in the Unrolling Model
@@
 5*  25* 26.96 16.58 18.07 12.17 5.27 11.78
5 20 27.52 17.36 18.97 12.51 5.30 11.99
5 30 27.25 17.27 18.77 12.51 5.29 11.96
4 25 27.84 17.23 17.96 12.43 5.32 11.93
6 25 26.97 16.92 17.97 12.46 5.25 11.80
@@ I.1 Optimization & Algorithm without DGTV Term
By removing the DGTV term $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ in Equation 4, and introducing the auxiliary variables ${\mathbf{z}}_{u},{\mathbf{z}}_{d}$ as Equation 9, we rewrite the unconstrained version of the optimization by augmented Lagrangian method as
$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{z}}_{d}+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$
$+({\boldsymbol{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|_{2}^{2}.$ (70)
@@
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}+{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau},$ (71)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (72)
$\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{z}}_{d}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}.$ (73)
The updating of ${\boldsymbol{\gamma}}_{u}^{\tau},{\boldsymbol{\gamma}}_{d}^{\tau}$ follows Equation 18.
@@ I.2 Optimization & Algorithm without DGLR Term
By removing the DGLR term ${\mathbf{x}}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ in Equation 4 and introduce only the auxiliary variables $\boldsymbol{\phi}$} in Equation 5 and ${\mathbf{z}}_{u}$ in Equation 9, we rewrite the unconstrained version of the optimization of ${\mathbf{x}}^{\tau+1}$ by augmented Lagrangian method as
$\min_{x,{\mathbf{z}}_{u}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{2}^{2}+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$ (74)
@@
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}$ $={\mathbf{H}}^{\top}{\mathbf{y}}+({\mathbf{L}}^{d}_{r})^{\top}(\frac{\gamma^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}$ (75)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}$ $=\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}$ (76)
The optimization of ${\boldsymbol{\phi}}^{\tau+1}$ is the same as Equation 15, and the solution is the same as Equation 16. The updating of ${\boldsymbol{\gamma}},{\boldsymbol{\gamma}}_{u}$ follows Equation 17 and Equation 18.